OpenStudy (anonymous):

As a fringe benefit you provide life insurance for your employees. What would a $132,000 policy cost if the charge was based on the following schedule. $10 per $1000 for the next $40,000 insurance $9 per $1000 for the next $40,000 insurance $7 per $1000 for any amount over $80,000

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

is it next then next or should it be first 40000 then next 40000 then after 80000

OpenStudy (anonymous):

40,000 then 40,000 then 80,000

OpenStudy (anonymous):

exactally how it looks above

OpenStudy (triciaal):

one approach 132000 -80000 = amount over 80000 divide by 1000 this gives # of 1000 over 80000 in 1000s cost $7 per 1000 for the 80000 -40000 = 40000 divide by 1000 to get # of 1000s cost $9 per 1000 next 40000 divide by 1000 to get #1000s cost = $10/1000 total cost over 80000 *7/1000 + 40*10 + 40*9 =

OpenStudy (triciaal):

they should use different numbers so you can really see the difference

OpenStudy (anonymous):

right!

OpenStudy (triciaal):

read this slow and let me know if you have any questions

OpenStudy (anonymous):

i dont get what formula to use on this

OpenStudy (triciaal):

did I specifically tell you to read slow?

OpenStudy (anonymous):

true ha

OpenStudy (anonymous):

sorry i didnt see the formula u gave me till now

OpenStudy (anonymous):

so i came up wih an answer of 1320

OpenStudy (anonymous):

now whats next?

OpenStudy (triciaal):

I got 1124

OpenStudy (triciaal):

cannot be right let me check

OpenStudy (anonymous):

no its right

OpenStudy (anonymous):

I have an answer book in the back of my book as well. your right

OpenStudy (anonymous):

I think i have to use PEMDAS

OpenStudy (triciaal):

the process is correct

OpenStudy (triciaal):

52*7 + 40*10+40*9= 364 +400 + 360 = 1124